Consider the ANOVA table that follows.

Analysis of Variance
Source DF SS MS F
Regression 4 4,105.03 1,026.26 23.75
Residual Error 49 2,117.05 43.21
Total 53 6,222.08
  1. a-1. Determine the standard error of estimate. (Round your answer to 2 decimal places.)

  1. a-2. About 95% of the residuals will be between what two values? (Round your answers to 2 decimal places.)

  1. b-1. Determine the coefficient of multiple determination. (Round your answer to 3 decimal places.)

  1. b-2. Determine the percentage variation for the independent variables. (Round your answer to 1 decimal place.)

  1. Determine the coefficient of multiple determination, adjusted for the degrees of freedom. (Round your answer to 3 decimal places.)
Answer #1

a1)standard error = sqrt(MSE)
= sqrt(43.21) = 6.57

a2)
0 +- 2 *SE
(-13.1469,13.1469)

b1)
R^2 = SS Regression / SS Total
= 4105.03 / 6222.08
= 0.6598


b2)
65.98 % of variation in y is explained by independ variables

adjusted R^2 = 1- (1- R^2)(n-1)/(n-k-1)
=
0.6320

Please give me a thumbs-up if this helps you out. Thank you! :)